OpenStudy (anonymous):

It takes Sally 4 hours to deliver newspapers. If Kathy works alone, it takes her 6 hours. How long would it take them working together to deliver all of the newspapers?

OpenStudy (anonymous):

(x/4)+(x/6)=1

OpenStudy (anonymous):

4x+6x=24 10x=24 x = 24/10 x =2.4 hours

OpenStudy (sriram):

x newspapers--- 4hrs x/4 newspapers in 1 hr x news papers--- 6 hrs x/6 news papers per hr let total time be t total news papers to be delivered again=x (x/4)*t+(x/6)*t=x dividing by x t/4+t/6=1

OpenStudy (anonymous):

click good answer if I desever